# Action accessible and weakly action representable varieties of algebras

### v1

@@ 1. Preliminaries
Let $\mathscr{C}$ be a semi-abelian category [11] and let $B,X$ be objects of a $\mathscr{C}$. We recall that a split extension of $B$ by $X$ is a diagram
@@
Examples of weakly action representable category include the variety $\mathbf{Assoc}$ of associative algebras [10], where $T=\operatorname{Bim}(X)$ is the associative algebra of bimultipliers of $X$ [17]; the variety $\mathbf{Leib}$ of Leibniz algebras, where $T=\operatorname{Bim}(X)$ is the Leibniz algebra of biderivations of $X$ [16, 18]; the varieties of $2$-nilpotent (commutative, anti-commutative and non-commutative) algebras (see [4, 13, 14, 15]). Note that in these mentioned examples, the weakly representing object is quite easy to consider. However, this is not always the case. The variety $\mathbf{CAssoc}$ of commutative associative algebras is also weakly action representable, but the algebra that could be expected to be a weakly representing object is not commutative. Nevertheless, in [4] a big colimit merged with the amalgamation property is used to naturally construct a weakly representing object.
@@
We recall that a span of monomorphisms $m\colon S\rightarrowtail B$ and $m^{\prime}\colon S\rightarrowtail B^{\prime}$ in a category $\mathscr{C}$ admits an amalgam in $\mathscr{C}$ if there exist two monomorphisms $u\colon B\rightarrowtail D$ and $u^{\prime}\colon B^{\prime}\rightarrowtail D^{\prime}$ in $\mathscr{C}$ such that $um=u^{\prime}m^{\prime}$. The relations between the representabilty of actions and the (AP) were firstly investigated in [2], where it was proved that for any Orzech category of interest, action representability is equivalent to the (AP) for protosplit monomorphisms, i.e., monomorphisms which form the kernel part of a split extension.
@@ 2. Nilpotent and $2$-solvable groups
$(b,x)(b^{\prime},x^{\prime})=(bb^{\prime},x\psi(g)(x^{\prime})),$
for any $b,b^{\prime}\in B$ and $x,x^{\prime}\in X$. Since $B$ is already $2$-nilpotent and $X$ is abelian, it is immediate to see that any nested commutators of length three must vanish. Thus, $B_{\psi}$ is a $3$-nilpotent group. The same result can be obtained for the semi-direct product $B_{\psi^{\prime}}$ and we can state the following.
@@ 3. Subvarieties of Lie algebras
Let $S=\mathbb{F}\{a,b\}$ be the $2$-dimensional abelian algebra and consider two copies $B,B^{\prime}$ of the $3$-dimensional Heisenberg algebra. More explicitly, $B=\mathbb{F}\{x,a,b\}$ with bracket $[x,a]=b$, and $B^{\prime}=\mathbb{F}\{y,a,b\}$ with bracket $[y,b]=a$, where the mononomorphisms are the obvious inclusions.
@@
Consider the Lie subalgebra $P$ of $D$ generated by $u(B)$ and $u^{\prime}(B^{\prime})$. Since $S$ is an ideal of $B$ and $B^{\prime}$, then $U=u(S)=u^{\prime}(S)$ is ideal of $P$. This gives us the adjoint map
@@
Hence, $\operatorname{ad}(P)$ is isomorphic to the special linear algebra $\mathfrak{sl}(2,\mathbb{F})$, which is a simple Lie algebra, and we get a contradiction. Thus, $m$ and $m^{\prime}$ cannot be amalgamated in any solvable Lie algebra. Since nilpotency implies solvability, this argument shows that the monomorphisms $m\colon S\rightarrowtail B$ and $m^{\prime}\colon X\rightarrowtail B^{\prime}$ cannot be amalgamated in any nilpotent Lie algebra. To conclude the proof, we observe that

### v2

@@ 1. Preliminaries
Let $\mathscr{C}$ be a semi-abelian category [11] and let $B,X$ be objects of $\mathscr{C}$. We recall that a split extension of $B$ by $X$ is a diagram
@@
Examples of weakly action representable category include the variety $\mathbf{Assoc}$ of associative algebras [10], where $T=\operatorname{Bim}(X)$ is the associative algebra of bimultipliers of $X$ [17]; the variety $\mathbf{Leib}$ of Leibniz algebras, where $T=\operatorname{Bider}(X)$ is the Leibniz algebra of biderivations of $X$ [16, 18]; the varieties of $2$-nilpotent (commutative, anti-commutative and non-commutative) algebras (see [4, 13, 14, 15]). Note that in these mentioned examples, the weakly representing object is quite easy to consider. However, this is not always the case. The variety $\mathbf{CAssoc}$ of commutative associative algebras is also weakly action representable, but the algebra that could be expected to be a weakly representing object is not commutative. Nevertheless, in [4] a big colimit merged with the amalgamation property is used to naturally construct a weakly representing object.
@@
We recall that a span of monomorphisms $m\colon S\rightarrowtail B$ and $m^{\prime}\colon S\rightarrowtail B^{\prime}$ in a category $\mathscr{C}$ admits an amalgam in $\mathscr{C}$ if there exist two monomorphisms $u\colon B\rightarrowtail D$ and $u^{\prime}\colon B^{\prime}\rightarrowtail D$ in $\mathscr{C}$ such that $um=u^{\prime}m^{\prime}$. The relations between the representabilty of actions and the (AP) were firstly investigated in [2], where it was proved that for any Orzech category of interest, action representability is equivalent to the (AP) for protosplit monomorphisms, i.e., monomorphisms which form the kernel part of a split extension.
@@ 2. Nilpotent and $2$-solvable groups
$(g,x)(g^{\prime},x^{\prime})=(gg^{\prime},x\psi(g)(x^{\prime})),$
for any $g,g^{\prime}\in B$ and $x,x^{\prime}\in X$. Since $B$ is $2$-nilpotent and $X$ is abelian, one may check that
$[B_{\psi},B_{\psi}]\cong\langle b\rangle\ltimes_{\psi}\mathbb{Z}_{5}^{2},$
$[B_{\psi},[B_{\psi},B_{\psi}]]\cong\mathbb{Z}_{5}$
and $[B_{\psi},[B_{\psi},[B_{\psi},B_{\psi}]]]$ is the trivial group. Thus, $B_{\psi}$ is a $3$-nilpotent group. The same result can be obtained for the semi-direct product $B_{\psi^{\prime}}$ and we can state the following.
@@ 3. Subvarieties of Lie algebras
Let $S=\mathbb{F}\{a,b\}$ be the $2$-dimensional abelian algebra and consider two copies $B,B^{\prime}$ of the $3$-dimensional Heisenberg algebra. More explicitly, $B=\mathbb{F}\{x,a,b\}$ with the only non-zero bracket $[x,a]=b$, and $B^{\prime}=\mathbb{F}\{y,a,b\}$ with the only non-zero bracket $[y,b]=a$, where the mononomorphisms are the obvious inclusions.
@@
Consider the Lie subalgebra $P$ of $D$ generated by $u(B)$ and $u^{\prime}(B^{\prime})$. Since $S$ is an ideal of $B$ and $B^{\prime}$, then $U=u(S)=u^{\prime}(S)$ is an ideal of $P$. This gives us the adjoint map
@@
Hence, $\operatorname{ad}(P)$ is isomorphic to the special linear algebra $\mathfrak{sl}(2,\mathbb{F})$, which is a simple Lie algebra, and we get a contradiction. Thus, $m$ and $m^{\prime}$ cannot be amalgamated in any solvable Lie algebra. Since nilpotency implies solvability, this argument shows that the monomorphisms $m\colon S\rightarrowtail B$ and $m^{\prime}\colon S\rightarrowtail B^{\prime}$ cannot be amalgamated in any nilpotent Lie algebra. To conclude the proof, we observe that
@@
Acknowledgements
The authors would like to thank the anonymous referees for their useful comments that helped us to improve the quality of the manuscript.